At the end of every quarter, Simon deposits $450 in an account that earns interest that is compounded quarterly. After 1.5 years, he will have $2874.48 in the account

How much total interest will have been earned?
Determine the annual rate of interest, compounded quarterly?

The answer is 174.48 and the annual rate of interest is 10% but how do u get it?
1
Expert's answer
2013-01-21T11:53:33-0500
Find how much money would be on Simon's account without percents.
450 * 18 = 2700
2874,48-2700 = 174.48 - his income for all the time.
So his cupping percentage is - 174.48 / 18 * 100 = 10%
ANSWER 174.48; 10%